OpenStudy (anonymous):

2+ 10x/x = x+5 what is LCD?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

since it is the only denominator you see

OpenStudy (anonymous):

2x+10=x^2+5x?

OpenStudy (phi):

notice that 10x/x is just 10. so simplify (always simplify if you can. It makes things simpler)

OpenStudy (ajprincess):



OpenStudy (anonymous):

I did that

OpenStudy (phi):

if you did, you would get 2+ 10 = x+5

OpenStudy (ajprincess):

u cud hav dun t b4 multiplyin evrythng by x.

OpenStudy (anonymous):

x(x+5)